(x^4D^4+6x^3D^3+9x^2D^2+3xD+1)y=(1+lnx)^2
1
Expert's answer
2021-07-23T05:42:44-0400

Given differential equation is:


"(x^4D^4+6x^3D^3+9x^2D^2+3xD+1)y=(1+lnx)^2"


Let us put, "x = e^z \implies z = ln(x), \frac{d}{dx} = D"


"(D(D-1)(D-2)(D-3)+6D(D-1)(D-2)+9D(D-1)+3D+1 )y = (1+z)^2"


"(D^4+2D^2+1)y = (1+z)^2"


The auxiliary equation will be,

"D^4+2D^2+1 =0"

"D = -i,-i,i,i"

Then,

"y = c_1(1+z)cos(z)+c_2(1+z)sin(z)"


P.I. "\frac{1}{D^4+2D^2+1}(1+z)^2 = \frac{1}{D^4+2D^2+1}(1+2z+z^2)"



"= \frac{1}{1+(D^4+2D^2)}(1+2z+z^2) = (1+(D^4+2D^2))^{-1}(1+2z+z^2)"


Expanding "(1+(D^4+2D^2))^{-1}" binomially.


"= (1-(D^4+2D^2)+\frac{-1(-1-1)}{2!}(D^4+2D^2)^2 + .........)(1+2z+z^2)"


"= (1+2z+z^2 - 2(2)) = (z^2+2z-3)"


The solution to the equation is,

"y = c_1(1+z)cos(z)+c_2(1+z)sin(z) + (z^2+2z-3)"



"y = c_1(1+lnx)cos(lnx) + c_2 (1+lnx)sin(lnx) + (lnx)^2+3lnx - 3"